Solve each equation for y to get two seperate functions that could be entered into a graphing calculator. Then combine both functions to create a single relation that involves x and y.
a. (y+2)squared=x
b. ysquared=x+2
c. (y+1)squared=x-6

Answer by stanbon(75887)

Solve each equation for y to get two seperate functions that could be entered into a graphing calculator. Then combine both functions to create a single relation that involves x and y.
a. (y+2)squared=x
---
Take the square root to get:
y+2 = Sqrt(x) or y+2 = -sqrt(x)
---
y = sqrt(x)-2 or y = -sqrt(x)-2
----------------------------------------
b. ysquared=x+2
Take the square root to get:
y = sqrt(x+2) or y = -sqrt(x+2)
Defined for All x >=(-2)
---------------------------------
c. (y+1)squared=x-6
Take the square root to get:
y+1 = sqrt(x-60)
y+1 = -sqrt(x-60)
Defined for All x >= 60
=========================
Cheers,
Stan H.